If Tommy bikes 2.55 x 10(negative 2 square) miles in 15 seconds, how many miles will he travel if he bikes consistantly for a week (no breaks)?
Your answer will be graded using the Delaware State Testing Rubric
(4,3,2,1, or 0).

Answer by checkley77(12844)

2.55*10^-2=.0255 Converting the number to a decimal.
.0255/15=.0017 MPS. Dividing by 15 to get miles for one by second.
or 60*60*.0017=6.12 MPH. Multiplying by 60 chnges seconds to minutes & multiplying by 60 again changes to hours thus converting mps to MPH.
6.12*24*7=1,028.16 MILES PER WEEK. Multiplying by 24 changes to days & multiplying by 7 changes to a week thus converting MPH to MPW.